OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y x - 5 y -x - 4

OpenStudy (anonymous):

no I did it the wrong way. When I write matrix I always write x on first column and y on second.

hero (hero):

actually, I agree, that 4.5 should be negative.

OpenStudy (anonymous):

now I really mess up haha don't listen to me

hero (hero):

y = x - 5 y = -x - 4 y = y x - 5 = -x - 4 x + x = 5 - 4 2x = 1 x = 0.5 y = -4.5

hero (hero):

Hey Romero, do you mind teaching me that method of solving using matrices?

OpenStudy (anonymous):

lol I'm just taking the class kinda hard to explain...

hero (hero):

It won't be difficult for me to understand. I'm a math guy, but let's go to vyew. It has a whiteboard, voice capabilities.

OpenStudy (anonymous):

get all variables on one side and all values to the left.line them up so the same variables line up in the same column remove the variables and try to get 1 on each row.

hero (hero):

I'm going to post a link to vyew in a few minutes.

hero (hero):

http://vyew.com/157292/Open_Study